# Hybrid data assimilation techniques using the adjoint method in a coupled Lorenz system

### v1

@@ Abstract
A hybrid 4D-variational data assimilation method for numerical climate models is introduced using the Lorenz ’63 model. This new approach has the potential to optimise a high complexity Earth system model (ESM) by utilising the adjoint equations of an intermediate complexity ESM. The method is conceptually demonstrated by consecutively synchronising two Lorenz ’63 systems to observations before optimisation. The first represents a ‘high complexity’ model and the second an ‘intermediate complexity’ model which has adjoint equations. This method will save computational power for a full ESM and has negligible error and uncertainty change compared to the optimisation of a single model with adjoint equations. A similar setup can be applied to sparse observations. An alternative assimilation setup, with two identical models, is used to filter noisy data. This reduces optimised parametric model uncertainty by approximately one third. Such a precision gain could prove valuable for seasonal, annual, and decadal predictions.
Keywords Lorenz ’63, data assimilation, adjoint, precision, uncertainty, climate models, error estimates
1 Introduction
@@
Due to the non-linearities within ESMs, the relevance of adjoint method, and thus the gradient of the cost function in reducing model-data misfits, can be limited by the predictability time scale. This can lead to spikes in the adjoint sensitivities which results in multiple local minima in the cost-function. Under such circumstances spikes occur in the estimated gradients and the cost function becomes very rough by showing an increasing number of local minima (Köhl and Willebrand, 2002; Lea et al., 2000). Fortunately the problem can be mitigated by synchronisation which removes the non-linear or chaotic dynamics from the adjoint model leading to a smooth cost function (Abarbanel et al., 2010; Sugiura et al., 2014). This method allows to extent the assimilation window beyond the predictability time-scale, provided that sufficient observations are available. Another challenge of applying an adjoint model to any state-of-the-art system is that these ESMs have a very large number of state variables $\mathcal{O}(10^{7}-10^{8})$, and the computation of its adjoint would be very time consuming and require large volumes of memory (Stammer et al., 2018).
@@
To mitigate both problems, we propose a novel framework where we use two climate models both coupled through synchronisation, one with a high resolution and complex parametrisation scheme and the other with coarse resolution and intermediate complexity but for which an adjoint exist. In this context we only use the adjoint of the low-resolution model to estimate the parameters that optimise the high-resolution model; i.e., that brings the model closer to the assimilated observations. A pre-requisite of this method is that the parameters being optimised must be the same in both models. The objective of this paper is to quantify the precision and the benefit of such a synchronised data assimilation approach. We perform this test conceptually using a Lorenz ’63 model system.
@@
The Lorenz ’63 system (Lorenz, 1963) is a well established proxy model to study chaotic fluid systems, such as the atmosphere (Gauthier, 1992; Miller et al., 1994; Pires et al., 1996; Stensrud and Bao, 1992; Kravtsov and Tsonis, 2021; Huai et al., 2017; Yang et al., 2006; Daron and Stainforth, 2015; Errico, 1997). The advantage is that it can be used to quantitatively evaluate the parameter dependence of the system prior to application in a full model. New modelling techniques can thus be trialled in relatively fast experiments (Pasini and Pelino, 2005; Tandeo et al., 2015; Goodliff et al., 2020; Marzban, 2013; Yin et al., 2014). It can also be used in a wide range of other applications (Du and Shiue, 2021; Cameron and Yang, 2019; Pelino and Maimone, 2007). The system generates a three dimensional, time varying trajectory which with variation of both model parameters and/or initial conditions will produce very different trajectories. Thus, it is an ideal test bed for testing non-linear modelling in a number of fields (Hirsch et al., 2013). In a previous study (Lyu et al., 2018) using the Lorenz ’63 model, the assimilation primarily focused on the fit of a single parameter $\rho$ and the initial conditions $(x,y,z)$ in a single model with an adjoint. This study will expand on this to fit all three model parameters simultaneously and use a model with adjoint to optimise the parameters of a model without one.
@@
The structure of the remaining paper is as follows: In Section 2 we introduce the model, outline the methodology of how these models are synchronised, show how the adjoint method is applied, describe our proposed multi-model setup, and detail the fitting procedure, including details on the minimisation algorithm. Section 3 shows and discusses the results of our multi-model setups, using a single model setup as a baseline for comparison. The results of introducing a mismodelling term to the adjoint model are also included along with a study of the setup’s response to sparse data conditions. Our results are summarised and conclusions discussed in Section 4.
@@ 2.2 Adjoint method
For the Lorenz ’63 system outlined in Eq. (2.1) the adjoint matrix can be derived by transposing its tangent liner matrix (TLM). The TLM follows from Eq. (2.1) as:
$M^{T}\equiv\frac{\partial\dot{\vec{x}}_{i}}{\partial\vec{x}_{j}}=\left(\begin{matrix}-\sigma&\sigma&0\\ \rho-z_{a}&-1&-x_{a}\\ y_{a}&x_{a}&-\beta\\ \end{matrix}\right).$ (2)
@@ 2.3 Synchronisation
A fundamental limitation of the adjoint method arises when integrating over periods that are longer than the predictability time scale of a system, leading to spikes in the estimated gradients and a cost function with an increasing number of local minima (Köhl and Willebrand, 2002; Lea et al., 2000). The problem can be mitigated by synchronisation which removes the non-linear or chaotic dynamics from the adjoint model (Abarbanel et al., 2010; Sugiura et al., 2014). To incorporate a synchronisation technique we expand the Lorenz ’63 model by so called nudging terms which then reads as:
@@
The adjoint matrix for the synchronised model in Eq. (2.3) is given by:
@@
There are seven possible combinations of the three state variables which can be synchronised. The effect of each of the possible choices on the root mean squared error (RMSE) between the true and adjoint systems is shown in Fig. 1 by varying the synchronisation constant $\alpha$ from 0 to 30. The figure demonstrates that synchronising the $z$-component is ineffective at reducing the RMSE (Yang et al., 2006). In contrast, synchronising both $x$ and $y$ prove effective, with $y$ leading the lowest RMSE values of the single variable for all values of $\alpha$. Synchronising $xyz$ and $xy$ achieve the most effective reduction in RMSE for the lowest value of $\alpha$. It can be seen that synchronising $z$ can lead to model instability. Thus, we choose to only synchronise $x$ and $y$ in the following, as seen in Eq. 2.3, to achieve more stable results with negligible precision loss.
@@
The Lorenz ’63 attractors for the trajectories of the true model and that with an adjoint are shown in Fig. 1(a) without synchronisation. A large divergence is visible between the attractors. However, if synchronisation is introduced, as shown in Fig. 1(b), the attractors become very similar. There is now significant overlap between their kernel density estimations (KDEs). We choose to use KDEs throughout this paper as they represent a smoothed estimate of the PDF for the attractor. This allows for convenient visual comparison of attractors. A more numerically rigorous method to check for effective synchronisation will be discussed in section 3.
(a) $\alpha=0$.
(b) $\alpha=10$.
Figure 2: The bottom left quadrants show the Lorenz ’63 truth and model attractors from the main three variable orientations. The diagonal plots show kernel density estimations (KDEs). Fig. 1(a) shows the attractors without synchronisation. Fig. 1(b) shows the attractors with synchronisation.
@@ 2.4 Multi-model data assimilation
This technique consecutively synchronises two forward models before running the adjoint of the second model backward in time. For this purpose, Eq. 2.3 must be modified to incorporate a consecutive synchronisation. A schematic of this setup is provided in Fig.3 and the implications of the two possible ways to calculate the cost function are discussed in the subsequent subsections. The equations of model 1, which is run only in forward mode, are
@@
To assimilate the data, we fit one of the synchronised models to the observations by optimising the model parameters. A cost function to calculate the misfit between observations and the model can be attached to model 1 or 2. The gradient of the cost function, with respect to the model parameters, is always calculated using the adjoint method associated with model 2. The adjoint model is numerically evaluated by automatic differentiation of the forward model 2. The process of synchronising all models, calculating the cost function and its gradient, and then adjusting model parameters is carried out iteratively by our chosen minimisation algorithm. Throughout all steps the parameter value of forward-only and adjoint models are identical and optimised simultaneously. The procedure stops when the algorithm-specific stopping criteria is met. The minimum of the cost function is the point at which the model provides the maximum likelihood estimate.
For the present study we have chosen to use the minimisation package iminuit (Dembinski and et al., 2020). The version currently used is the Minuit2 algorithm MIGRAD which is a variable-metric method including an inexact line search, a stable metric updating scheme, and it checks for positive-definiteness. This is chosen as it is a robust and stable minimiser that both utilises the gradient we derived using adjoint model in Eq. (11), and calculates accurate uncertainty estimates for all fitted parameters. Further details on the specific minimisation algorithm are given in James and Roos (1975).
@@ 2.4.1 Setup 1 - state filtered data assimilation (SFDA)
Assuming that both Lorenz models can be thought of as representing two identical climate models, the cost function can be placed on the model 2 which has adjoint equations. This allows model 1 to filter out some of the background noise on the observations before they are given to the cost function attached to model 2. We will subsequently refer to this setup as state filtered data assimilation (SFDA.) This filtering should act to increase the sensitivity of the cost function to the parameters by reducing the impact of observational noise which favourably increases the signal to background ratio synchronised into model 2. Such a filtering setup would theoretically reduce parametric uncertainty below that of traditional single model data assimilation because model 1 should act to reduce the amount of noise synchronised into model 2. This produces a steeper, more-pronounced cost function increasing the parametric sensitivity which reduces the uncertainty on each fitted parameter.
In SFDA the cost function acts to constrain model 2 which has an adjoint. The cost function is
@@
$N$ is the total number of time steps of the assimilation window, and $\sigma_{\vec{x}_{o}}$ is the uncertainty associated with the observation noise. The gradient will be calculated subsequently using the adjoint of model 2.
@@ 2.4.2 Setup 2 - hybrid data assimilation (HDA)
Here we want to explore if using an existing adjoint from one model could be utilised to optimise a different target model without adjoint. This will be referred to as hybrid data assimilation (HDA.) In HDA we assume that both models differ in their equations and spatial resolution. Instead of interpolating or transforming the original model onto the adjoint model grid, formulation of the adjoint model through synchronisation would provide a simpler means to do this, as only essential parameters need to be optimised. Auxiliary variables and parameters, including those that may not exist in the target model, will be adjusted via the model dynamics.
@@
$N$ is the total number of time steps of the assimilation window and $\sigma_{\vec{x}_{o}}$ is the uncertainty associated with the observation noise. This measures the quadratic misfit between the forward-only model 1 and the observations. Model 1 $\vec{x}_{f}$ will be constrained by this cost function and its gradient will be calculated using the adjoint of model 2 $\vec{x}_{a}$.
@@ 2.4.3 Cost function gradient
The gradient of the cost function, with respect to the parameters, will be identical for both methods above. This is because adjoint equations will always be associated with model 2. The gradient of the cost function, with respect to the state variables at $t=0$, is given by
$\vec{\nabla}_{\vec{x}_{a}}J=\frac{1}{N}\vec{\lambda}_{\vec{x}_{a}}(0),$ (9)
where $\vec{\lambda}$ is the adjoint vector with respect to the state variables calculated in the reverse time direction:
$\vec{\lambda}_{\vec{x}_{a}}(N+1)$ $=$ $\vec{0},$ (10a)
$\vec{\lambda}_{\vec{x}_{a}}(t)$ $=$ $\vec{x}_{o}^{\ast}(t)\frac{1}{\sigma_{\vec{x}_{o}}}\left(\vec{x}_{o}(t)-\vec{x}_{a}(t)\right)$
$-M^{\ast}(t)\vec{\lambda}_{\vec{x}_{a}}(t+1).~{}\text{for}~{}t=N,...,0$
$M^{\ast}$ is the adjoint matrix, calculated by automatic differentiation, and $\vec{x}_{o}^{\ast}(t)$ is the adjoint of the observation. These equations were derived using the method detailed in Talagrand (2010). This gradient can then be calculated with respect to the parameters $(\sigma,\rho,\beta)$ notated by the subscript $\vec{p}$. This yields
$\vec{\nabla}_{\vec{p}}J=\frac{1}{N}\sum_{t=N}^{0}\vec{\lambda}_{\vec{x}_{a}}(t)\cdot\left(\begin{matrix}y_{a}(t)-x_{a}(t)\\ x_{a}(t)\\ -z_{a}(t)\end{matrix}\right)$ (11)
where the vector derived using the chain rule is multiplied by each term from Eq. (2.4.3) before being summed.

### v3

# 1]Universität Hamburg, Fakultät für Mathematik, Informatik und Naturwissenschaften, Fernerkundung & Assimilation, Bundesstr. 53, 20146 Hamburg, Deutschland

@@ Abstract
A hybrid 4D-variational data assimilation method for chaotic climate models is introduced using the Lorenz ’63 model. This approach aims to optimise an Earth system model (ESM), for which no adjoint exists, by utilising an adjoint model of a different, potentially simpler ESM. The technique relies on synchronisation of the model to observed time series data employing the dynamical state and parameter estimation (DSPE) method to stabilise the tangent linear system by reducing all positive Lyapunov exponents to negative values. Therefore, long windows can be used to improve parameter estimation. In this new extension a second layer of synchronisation is added between the two models, with and without an adjoint, to facilitate linearisation around the trajectory of the model without an adjoint. The method is conceptually demonstrated by synchronising two Lorenz ’63 systems, representing two ESMs, one with and the other without an adjoint model. Results are presented for an idealised case of identical, perfect models and for a more realistic case in which they differ from one another. If employed with a coarser ESM with an adjoint, the method will save computational power as only one forward run with the full ESM per iteration needs to be carried out. It is demonstrated that there is negligible error and uncertainty change compared to the ‘traditional’ optimisation of full ESM with an adjoint. In a variation of the method outlined, synchronisation between two identical models can be used to filter noisy data. This reduces optimised parametric model uncertainty by approximately one third. Such a precision gain could prove valuable for seasonal, annual, and decadal predictions.
[…]
@@
Due to the non-linearities within ESMs, the relevance of adjoint method, and thus the gradient of the cost function in reducing model-data misfits, can be limited by the predictability time scale. This can lead to spikes in the adjoint sensitivities as a result of multiple local minima in the cost-function. Under such circumstances spikes occur in the estimated gradients and the cost function becomes very rough by showing an increasing number of local minima (Köhl and Willebrand, 2002; Lea et al., 2000). Fortunately, the problem can be mitigated by synchronisation which removes the non-linear or chaotic dynamics from the adjoint model leading to a smooth cost function (Abarbanel et al., 2010; Sugiura et al., 2014). This method allows extension of the assimilation window beyond the predictability time-scale, provided that sufficient observations are available. Another challenge of applying an adjoint model to any state-of-the-art system is that these ESMs have a very large number of state variables $\mathcal{O}(10^{7}-10^{8})$, and the computation of its adjoint would be very time consuming and require significant volumes of memory (Stammer et al., 2018).
@@
To mitigate both problems, we propose a novel framework where we use two climate models both coupled through synchronisation, one with a high resolution and the other with coarse resolution for which an adjoint exists. In this context we only use the adjoint of the low-resolution model to estimate the parameters that optimise the high-resolution model; i.e., that brings the model closer to the assimilated observations. A pre-requisite of this method is that the parameters being optimised must be the same in both models. The objective of this paper is to quantify the precision and the benefit of such a synchronised data assimilation approach. We perform this test conceptually using a Lorenz ’63 model system.
@@
The Lorenz ’63 system (Lorenz, 1963) is a well established proxy model to study chaotic fluid systems, such as the atmosphere (Gauthier, 1992; Miller et al., 1994; Pires et al., 1996; Stensrud and Bao, 1992; Kravtsov and Tsonis, 2021; Huai et al., 2017; Yang et al., 2006; Daron and Stainforth, 2015; Errico, 1997). The advantage is that it can be used to quantitatively evaluate the parameter dependence of the system prior to application in a full model. New modelling techniques can thus be trialled in relatively fast experiments (Pasini and Pelino, 2005; Tandeo et al., 2015; Goodliff et al., 2020; Marzban, 2013; Yin et al., 2014). It can also be used in a wide range of other applications (Du and Shiue, 2021; Cameron and Yang, 2019; Pelino and Maimone, 2007). The system generates a three dimensional, time varying trajectory which with variation of both model parameters and/or initial conditions will produce very different trajectories. Thus, it is an ideal test bed for non-linear modelling in a number of fields (Hirsch et al., 2013). In a previous study (Lyu et al., 2018) using the Lorenz ’63 model, the assimilation primarily focused on the fit of a single parameter $\rho$ and the initial conditions $(x,y,z)$ in a single model with an adjoint. This study will expand on the previous to fit all three model parameters simultaneously and use a model with adjoint to optimise the parameters of a model without one.
@@
The structure of the remaining paper is as follows: In Section 1 we introduce the model, outline the methodology of how these models are synchronised, show how the adjoint method is applied, describe our proposed multi-model setup, and detail the fitting procedure, including the minimisation algorithm. Section 2 shows and discusses the results of our multi-model setups, using a single model setup as a baseline for comparison. The results of introducing a mismodelling term to the adjoint model are also included. Our results are summarised and conclusions discussed in Section 2.3.
@@ 1.2 Adjoint method
For the Lorenz ’63 system outlined in Eq. (1) the adjoint matrix can be derived by transposing its tangent liner matrix (TLM). The TLM follows from Eq. (1) as:
$M^{T}\equiv\frac{\partial\dot{\vec{x}}_{i}}{\partial\vec{x}_{j}}=\left(\begin{matrix}-\sigma&\sigma&0\\ \rho-z&-1&-x\\ y&x&-\beta\\ \end{matrix}\right).$ (2)
@@ 1.3 Synchronisation
A fundamental limitation of the adjoint method arises when integrating over periods that are longer than the predictability time scale of a system, leading to exponentially growing gradients and a cost function with an increasing number of local minima (Köhl and Willebrand, 2002; Lea et al., 2000). The problem can be mitigated by synchronisation which removes the non-linear or chaotic dynamics which effect the cost function (Abarbanel et al., 2010; Sugiura et al., 2014). To incorporate a synchronisation technique, we expand the Lorenz ’63 model by adding so-called nudging terms which then reads:
@@
The adjoint matrix for the synchronised model in Eq. (3) is given by:
@@
There are seven possible combinations of the three state variables which can be synchronised. The effect of each of the possible choices on the root mean squared error (RMSE) between the true and adjoint systems is shown in Fig. 1 by varying the synchronisation constant $\alpha$ from 0 to 30. The figure demonstrates that synchronising the $z$-component is ineffective at reducing the RMSE (Yang et al., 2006). In contrast, synchronising both $x$ and $y$ prove effective, with $y$ leading the lowest RMSE values of the single variable for all values of $\alpha$. Synchronising $xyz$ and $xy$ achieve the most effective reduction in RMSE for the lowest value of $\alpha$. It can be seen that synchronising $z$ can lead to model instability. Thus, we choose to only synchronise $x$ and $y$ in the following, as seen in Eq. 3, to achieve more stable results with negligible precision loss.
@@
The Lorenz ’63 attractors for the trajectories of the true model and that with an adjoint are shown in Fig. 2(a) without synchronisation. A large divergence is visible between the trajectories. However, if synchronisation is introduced, as shown in Fig. 2(b), the trajectories become very similar. There is now significant overlap between their kernel density estimations (KDEs). We choose to use KDEs throughout this paper as they represent a smoothed estimate of the PDF for the trajectory. This allows for convenient visual comparison of trajectories. A more numerically rigorous method to check for effective synchronisation will be discussed in section 2.
((a)) $\alpha=0$.
((b)) $\alpha=10$.
[…]
Figure 2: The bottom left quadrants show the Lorenz ’63 truth and model attractors from the main three variable orientations. The diagonal plots show kernel density estimations (KDEs). Fig. 2(a) shows the trajectories without synchronisation. Fig. 2(b) shows the trajectories with synchronisation.
@@ 1.4 Multi-model data assimilation
This technique consecutively synchronises two forward models before running the adjoint of the second model backward in time. For this purpose, Eq. 3 must be modified to incorporate a consecutive synchronisation. A schematic of this setup is provided in Fig.3 and the implications of the two possible ways to calculate the cost function are discussed in the subsequent subsections. The equations of model 1, which is run only in forward mode, are
@@
To assimilate the data, we fit one of the synchronised models to the observations by optimising the model parameters. A cost function is constructed to calculate the misfit between observations and the model of interest. The gradient of the cost function, with respect to the model parameters, is always calculated using the adjoint method associated with model 2. However, the form of the adjoint will vary between the two methods we subsequently present. The adjoint model is numerically evaluated by automatic differentiation (AD) of the forward model 2 with respect to either model 1 or 2 depending on our chosen setup. This is done in the python package JAX which numerically evaluates the vector Jacobian product of the model with respect to its state variable vector (Bradbury et al., 2018). This is then integrated using an inverse Runge-Kutta scheme. Our code stores the state variables and adjoint vectors at each time step. It is also possible to carry out the entire integration using JAX. The process of synchronising all models, calculating the cost function and its gradient, and then adjusting model parameters is carried out iteratively by our chosen minimisation algorithm. Throughout all steps the parameter value of forward-only and adjoint models are identical and optimised simultaneously.
[…]
@@ 1.4.1 Setup 1 - state filtered data assimilation (SFDA)
Assuming that both Lorenz models can be thought of as representing two identical climate models, the cost function can be placed on the model 2. This allows model 1 to filter out some of the background noise on the observations before they are given to the cost function attached to model 2. This favourably increases the signal to background ratio synchronised into model 2. Such a filtering setup would theoretically reduce parametric uncertainty below that of traditional single model data assimilation because model 1 should act to reduce the amount of noise synchronised into model 2. We will subsequently refer to this setup as state filtered data assimilation (SFDA.)
In SFDA the cost function acts to constrain model 2. The cost function is
@@
$N$ is the total number of time steps of the assimilation window, and $\sigma_{\vec{x}_{o}}$ is the uncertainty associated with the observation noise. The adjoint matrix includes additional terms, compared with Eq. 4, arising from the second synchronisation step for model 2. This is given by:
$M_{\text{SFDA}}^{\ast}=\left(\begin{matrix}-(\sigma+\alpha)&\rho-z_{a}&y_{a}&0&0&0\\ \sigma&-(1+\alpha)&x_{a}&0&0&0\\ 0&-x_{a}&-\beta&0&0&0\\ \alpha&0&0&-(\sigma+\alpha)&\rho-z_{f}&y_{f}\\ 0&\alpha&0&\sigma&-(1+\alpha)&x_{f}\\ 0&0&0&0&-x_{f}&-\beta\\ \end{matrix}\right)$ (8)
which in practice is numerically evaluated using AD.
@@ 1.4.2 Setup 2 - hybrid data assimilation (HDA)
Here we want to explore if using an existing adjoint from one model could be utilised to optimise a different target model without adjoint. This will be referred to as hybrid data assimilation (HDA.) In HDA we assume that both models may differ in resolution or numerical formulation but are governed by the same equations. Instead of interpolating or transforming the original model variables onto the adjoint model grid, formulation of the adjoint model through synchronisation would provide a simpler means to do this, as only essential parameters need to be interpolated. Auxiliary variables and parameters, including those that may not exist in the target model, will be generated by the synchronised model.
@@
$N$ is the total number of time steps of the assimilation window and $\sigma_{\vec{x}_{o}}$ is the uncertainty associated with the observation noise. This measures the quadratic misfit between the forward-only model 1 and the observations. Model 1 $\vec{x}_{f}$ will be constrained by this cost function and its gradient will be calculated using the adjoint of model 2 $\vec{x}_{a}$. The adjoint matrix is
$M_{\text{HDA}}^{\ast}=\left(\begin{matrix}-(\sigma+\alpha)&\rho-z_{a}&y_{a}\\ \sigma&-(1+\alpha)&x_{a}\\ 0&-x_{a}&-\beta\\ \end{matrix}\right).$ (10)
which is numerically evaluated using AD. The synchronisation with model 1 ensures that the trajectory of model 2 $\vec{x}_{a}$ closely follows that of model 1.
@@ 1.4.3 Cost function gradient
The gradient of the cost function, with respect to the state variables at $t=0$, is given by
$\vec{\nabla}_{\vec{x}_{a}}J=\frac{1}{N}\vec{\lambda}(0),$ (11)
where $\vec{\lambda}$ is the adjoint vector. The adjoint vector is calculated by integration of the differential adjoint equation in the reverse time direction. This will differ between our two setups.
SFDA (setup 1)
This adjoint equation for SFDA is given by:
$\dot{\vec{\lambda}}_{\text{SFDA}}(t)$ $=$ $\frac{1}{\sigma_{\vec{x}_{o}}^{2}}\left(\vec{x}_{o}(t)-\vec{x}_{a}(t)\right)$
$-M_{\text{SFDA}}^{\ast}(t)\vec{\lambda}_{\text{SFDA}}(t)\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ t=N,...,0$
$\text{with}\leavevmode\nobreak\ \leavevmode\nobreak\ \vec{\lambda}_{\text{SFDA}}(N)$ $=$ $\vec{0}.$ (12b)
These equations were derived using the method detailed in Talagrand (2010). The gradient can then be calculated with respect to the parameters $(\sigma,\rho,\beta)$ notated by the subscript $\vec{\theta}$. This yields
$\vec{\nabla}_{\vec{\theta}}J_{\text{SFDA}}=\frac{1}{N}\sum_{t=N}^{0}\vec{\lambda}_{\text{SFDA}}(t)\left(\begin{matrix}y_{a}(t)-x_{a}(t)\\ x_{a}(t)\\ -z_{a}(t)\\ y_{f}(t)-x_{f}(t)\\ x_{f}(t)\\ -z_{f}(t)\end{matrix}\right)$ (13)
which is a component-wise multiplication at each time step.
HDA (setup 2)
This adjoint equation for HDA is given by:
$\dot{\vec{\lambda}}_{\text{HDA}}(t)$ $=$ $\frac{1}{\sigma_{\vec{x}_{o}}^{2}}\left(\vec{x}_{o}(t)-\vec{x}_{f}(t)\right)$
$-M_{\text{HDA}}^{\ast}(t)\vec{\lambda}_{\text{HDA}}(t)\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ t=N,...,0$
$\text{with}\leavevmode\nobreak\ \leavevmode\nobreak\ \vec{\lambda}_{\text{HDA}}(N)$ $=$ $\vec{0}.$ (14b)
The gradient with respect to the parameters $\vec{\theta}=(\sigma,\rho,\beta)$ is calculated to be:
$\vec{\nabla}_{\vec{\theta}}J_{\text{HDA}}=\frac{1}{N}\sum_{t=N}^{0}\vec{\lambda}_{\text{HDA}}(t)\left(\begin{matrix}y_{a}(t)-x_{a}(t)\\ x_{a}(t)\\ -z_{a}(t)\end{matrix}\right)$ (15)
which is again a component-wise multiplication at each time step. For HDA and SFDA, the trajectories of both models and adjoint vectors are stored for evaluation of the gradient.